# News is More than a Collection of Facts: Moral Frame Preserving News Summarization

### v1

@@ 1 Introduction
With shrinking attention spans in the social media era, fewer people read full news articles, favoring concise information, instead (Baqir et al., 2024). News summarization addresses this shift by condensing content into digestible formats. AI-driven summarization has long been studied (Zhang et al., 2024a), with algorithms traditionally evaluated on fluency and coherence of the generated text, and faithfulness to the facts presented in the article (Fabbri et al., 2021). However, journalism researchers recognize that news articles are more than just a collection of facts—they contain intentional framing (De Vreese, 2005).
@@
We perform the first exploration into the preservation of moral framing in AI-generated summaries. Precisely, we aim to identify a prompting strategy that best preserves moral framing but maintains overall summary quality. We leverage the zero-shot summarization ability of Large Language Models, shown to produce results on par with human-generated summaries (Goyal et al., 2022; Zhang et al., 2024b). We compare three language models and five prompting methods. Leveraging the intuition that journalists intentionally use or report moral-laden words in the article text, we propose approaches that first identify moral-laden words in the article (e.g., through Chain-of-Thought or supervised classification) and then guide the language model in preserving such words in the summary.
@@
Figure 1: Example of moral framing, sourced from a 2002 Kyodo News article in the EMONA dataset (Lei et al., 2024). On the left, an example of how an AI-generated summary might overlook the moral framing in this segment. On the right, an example of how our proposed approach identifies the moral-laden words and preserves (part of) them in the summary.
@@ 2.1 Automated text summarization
AI-based summarization has a long history (Zhang et al., 2024a), initially relying on extractive summarization techniques that identify the most significant spans of text to include in the summary (Erkan & Radev, 2004; Mihalcea & Tarau, 2004). The development of deep learning and transformers shifted the focus to abstractive summarization, which generates summaries from scratch by fine-tuning pre-trained models, e.g., BERT (Liu & Lapata, 2019) or BART (Liu et al., 2022), on reference summaries. The advent of Large Language Models (LLMs) shifted the attention away from the fine-tuning paradigm. Large benchmarking studies on news summarization have shown that humans judge summaries generated by LLMs through zero-shot prompting to be comparable to human-written ones, and overwhelmingly prefer them over the reference summaries traditionally used to fine-tune language models (Goyal et al., 2022; Pu et al., 2023; Zhang et al., 2024b). Recent work has focused on the refinement of the summarization prompts, e.g., through Chain-of-Thought approaches that first prompt the model to list important facts and then integrate these facts into a summary (Wang et al., 2023) or that guide the model in retrieving the relevant context (Liu et al., 2024).
@@ 4 Experiments
We describe our experimental procedure starting with the dataset and models used.
@@ 4.1 Dataset
We experiment with the EMONA dataset (Lei et al., 2024), a collection of 400 news articles spanning different topics and political biases. In each sentence of the articles, all event instances were annotated by humans—where an event refers to an occurrence or action, considered the basic element in storytelling (Zhang et al., 2021)—yielding a total of over 10k sentences and 45k events. Each event was then labeled as moral-laden or neutral, with 21% (9.6k) of the events labeled as moral-laden. When labeled as moral-laden, the event was then annotated with one of the ten moral elements of the Moral Foundation Theory.
@@ 4.2 Models
We compare summarization models that demonstrate strong general knowledge understanding and are instruction-tuned, as recommended by Zhang et al. (2024b). Specifically, we compare Llama-3-70B, Command R+, and DeepSeek R1, three state-of-the-art open-source instruction-tuned models. For the Class method, we train Llama-3-8B to identify moral-laden words in the article text in a supervised fashion using the EMONA dataset (we employ a smaller model than the summarization models to keep training and validation manageable). We feed the article’s sentences individually to the classifier which predicts moral-ladenness at a token level. We also tested two other approaches, namely (1) using the full article as input, and (2) a sequential approach that first identifies the sentences that contain moral-laden words and then the moral-laden words in them, but found them to perform worse. Hyperparameters and additional details are reported in Appendix B.
@@ 4.3 Summaries generation procedure
To obtain the classifier for the Class method, we (1) perform a 3-fold cross-validation dividing the training set further into training and validation sets and choose the hyperparameters that lead to the best performance on the validation set, (2) train the model on the whole training set, and (3) perform inference on the test set’s articles. We then generate summaries of the test set articles with the five methods and three summarization models. To account for variability, we repeat the generation five times with fixed seeds and report the average results. We randomly choose a seed whose results are used for human evaluation. For the word-preserving approaches, we treat all annotated moral-laden events as human-annotated moral-laden words, and all other article words as neutral. We do not distinguish among the ten MFT elements in the summary generation, but we do in the evaluation.
@@ 4.4 Evaluation
We perform automated, crowd, and expert evaluations of the summaries. Our evaluations are not intended to judge the overall quality of LLM-generated summaries, which has already been extensively studied (Goyal et al., 2022; Pu et al., 2023; Zhang et al., 2024b). Instead, we employ these evaluations to compare moral framing preservation across methods and models, and assess whether instructing the summarizer to preserve moral-laden words compromises the overall quality of the summary. First, we employ five automated evaluation metrics to compare the generated summaries across summarization models. We use these results to choose the best-performing summarization model. Then, we perform crowd and expert evaluations of the summaries generated by the chosen model on the test set to investigate the aspects of moral framing preservation that are not captured by the automated metrics. Appendix C provides additional details on the human evaluations.
@@ 4.4.1 Automated evaluation
In line with a recent benchmarking survey (Zhang et al., 2024b), we employ the following reference-free automated evaluation metrics to measure the quality of the summaries: (1) SummaC(Laban et al., 2022), which measures how the summary sentences are entailed from the news article; (2) QAFactEval(Fabbri et al., 2022), which measures entailment based on questions generated from the article and answers based on the summary; and (3) BLANC(Vasilyev et al., 2020), which measures the performance boost gained by a language model with access to the summary while performing a language understanding task on the article.
In addition, we introduce the following metrics to measure the preservation of the moral-laden words in the generated summaries: (1) Moral Count (MC), which reports the number of words that were annotated as moral-laden in the article and that are preserved in the summary (to avoid double-counting, we employ the lemmas of the annotated and preserved words). (2) Moral Divergence (MD), which measures the divergence between the distribution of moral annotations in the article and the summary. We normalize the distribution of moral annotations across the ten MFT moral categories and employ the Jensen-Shannon divergence to calculate the divergence between the distribution in the article and the summary (smaller divergence indicates more similarity between the distributions).
@@ 5.1 Automated evaluation
Table 2: Results of the automated evaluation of the summaries of the test set articles.
$\uparrow$ indicates that higher scores are better, $\downarrow$ indicates the opposite.
Second, on MC and MD, the five methods rank in the order: Plain, Direct, CoT, Class, Oracle, consistently across the three models. Thus, instructing the summarizer to preserve moral framing even without providing a list of moral-laden words (Direct) leads to better preservation of moral-laden words than just instructing to generate a summary (Plain).
@@
As expected, the word-preserving methods outperform the basic methods on MC and MD. To compare the three methods, we first report the word-level moral-ladenness prediction performance. Recall (Table 1) that, in CoT, the summarizer was prompted to identify the moral-laden words to be preserved in the summary; in Class, a classifier was trained to detect moral-laden words in the article. The efficacy of these components is instrumental in the subsequent summary generation with the two methods. Table 3 reports their performance, measured by comparing the lists of predicted words and the list of words annotated as moral-laden in the EMONA dataset (thus, random prediction would lead to a near-zero $F_{1}$-score). We observe that Class leads to better moral-ladenness prediction results than CoT, which is reflected in MC and MD results in Table 2. Finally, Oracle—where the summarizer is provided with the list of words annotated as moral-laden—unsurprisingly, outperforms the other two word-preserving methods.
@@ 5.2 Crowd evaluation
Table 4 reports, for each method, the average Likert scores the workers gave to the morally framed spans they highlighted, and the differences between the scores assigned to the summaries generated by different methods. First, we observe that Class and Oracle emerge as the best-performing methods with significant differences from the other three methods. However, we notice no significant difference between the two although Oracle is more effective in preserving the annotated moral-laden words than Class, which in turn is more effective than CoT (as shown in Section 5.1). These results suggest that, when a sufficient moral-ladenness prediction performance is met, preserving moral-laden words in the summary improves the preservation of moral framing, but a better moral-ladenness prediction performance does not necessarily correlate with better moral framing preservation.
@@
Next, we observe that all average scores range between 3 and 3.5. This is not surprising since these scores evaluate the extent to which morally framed article spans are preserved in the summary, and some are inevitably lost due to the summary’s shorter length.
Finally, we find no significant correlations between automated and crowd evaluation results (as reported in Table 5.2 together with the correlation to the summaries length). Thus, judging moral framing preservation requires human evaluation, as it (1) is not correlated to the summary quality, and (2) cannot be assessed by measuring the preservation of moral words.
@@
To evaluate the effect of preserving moral-laden words in the summary, we examine the scores given to the highlighted spans containing at least one annotated moral-laden word that is preserved in the summary. The spans containing moral-laden words are the same for all summaries—however, some summaries preserve these words while others do not. Thus, the number of these spans varies across methods. Table 6 reports the average scores assigned to such spans and their number. We observe that the scores are similar across methods and consistently higher than the average scores presented in Table 4, suggesting that maintaining moral-laden words promotes the preservation of moral framing. The difference across methods lies in the number of such highlighted spans, with Class and Oracle counting most of such spans (respectively, 24% and 39% more when compared to Plain), leading to the difference in the average scores observed in Table 4.
@@
The workers were allowed to highlight as many morally framed spans as necessary, resulting in a number of highlights ranging from 0 (in a single case) to 31, with an average of $6.6\pm 4.7$. We correlate the number of spans a worker highlighted to the Likert scores they assigned to the summaries. We group the number of highlights per article in blocks of three and report in Figure 2 the average scores assigned by the workers to each method’s summaries. The scores exhibit a downward trend—the more spans a worker highlighted, the lower the average scores they assigned to them. This is expected since the more text spans are highlighted, the more challenging it is for the summarizer to preserve them all in the summary. Next, we observe that Class generally outperforms the other methods across the numbers of highlights. In contrast, Oracle leads to the worst results in the 1-2-3 range but the best in the 4-5-6 range. We conjecture that this is due to the number of moral-laden words to be preserved in the summary (which is the highest for Oracle, as reported in Appendix D.1)—attempting to preserve a large number of words may dilute the reporting of key aspects of the article, which is especially noticeable when only a few key morally framed text spans are highlighted and evaluated.
1-2-3 (23%)4-5-6 (40%)7-8-9 (20%)10-11-12 (9%$\geq$13 (8%)222233334444Average scorePlainDirectCoTOracleClass
@@ 6 Conclusions
Explaining how the summarization models identify moral-laden words and decide which ones to preserve is instrumental to improving transparency to users. Further, reducing the reliance on human annotations can foster the approach’s practicality, e.g., by enhancing the Chain-of-Thought method with few-shot examples. Finally, the EMONA dataset offers further avenues of exploration, e.g., our results could be correlated with the political bias of the article’s source, and the prompting methods could be extended by integrating the Moral Foundation Theory annotations associated with moral-laden words.
@@
We provide all the necessary details for reproducing our experiments throughout the paper and the Appendix. Methods, prompting, and training procedures are described in Sections 3 and 4. Appendix A reports the complete prompts used to generate the summaries. Appendix B provides additional details on our experiments, including the tested and used hyperparameters and computational infrastructure. Appendix C presents additional information on the human evaluations, including the aggregated demographics and further details on the crowd task. The code, generated summaries, and (anonymized) automated, crowd, and expert evaluation results are provided as supplemental material. The model weights trained with the Class method will be released at publication due to size limitations.
Acknowledgments
This material is produced as part of AlgoSoc, a collaborative 10-year research program on public values in the algorithmic society, and the Hybrid Intelligence Centre, both funded by the Dutch Ministry of Education, Culture and Science (OCW) under the Gravitation programme (project numbers 024.005.017 and 024.004.022). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of OCW or those of the AlgoSoc consortium as a whole. Michela Lorandi’s work was conducted with the financial support of the Science Foundation Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real) under Grant No 18/CRT/6224 and the ADAPT SFI Research Centre at Dublin City University, funded by the Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2.

### v2

@@ 1 Introduction
With shrinking attention spans in the social media era, fewer people read full news articles, favoring concise information, instead (Baqir et al., 2024). News summarization addresses this shift by condensing content into digestible formats. AI-driven summarization has long been studied (Zhang et al., 2025), with algorithms traditionally evaluated on fluency and coherence of the generated text, and faithfulness to the facts presented in the article (Fabbri et al., 2021). However, journalism researchers recognize that news articles are more than just a collection of facts—they contain intentional framing (De Vreese, 2005).
@@
We perform the first exploration into the preservation of moral framing in AI-generated summaries. Precisely, we aim to identify a prompting strategy that best preserves moral framing but maintains overall summary quality. We leverage the zero-shot summarization ability of Large Language Models, shown to produce results on par with human-generated summaries (Goyal et al., 2022; Zhang et al., 2024). We compare three language models and five prompting methods. Leveraging the intuition that journalists intentionally use or report moral-laden words in the article text, we propose approaches that first identify moral-laden words in the article (e.g., through Chain-of-Thought or supervised classification) and then guide the language model in preserving such words in the summary.
@@
Figure 1: Example of moral framing, sourced from a 2002 Kyodo News article in the EMONA dataset (Lei et al., 2024). At the top, an example of how an AI-generated summary might overlook the moral framing in this segment. At the bottom, an example of how our proposed approach identifies the moral-laden words and preserves (part of) them in the summary.
@@ 2.1 Automated text summarization
AI-based summarization has a long history (Zhang et al., 2025), initially relying on extractive summarization techniques that identify the most significant spans of text to include in the summary (Erkan & Radev, 2004; Mihalcea & Tarau, 2004). The development of deep learning and transformers shifted the focus to abstractive summarization, which generates summaries from scratch by fine-tuning pre-trained models, e.g., BERT (Liu & Lapata, 2019) or BART (Liu et al., 2022), on reference summaries. The advent of Large Language Models (LLMs) shifted the attention away from the fine-tuning paradigm. Large benchmarking studies on news summarization have shown that humans judge summaries generated by LLMs through zero-shot prompting to be comparable to human-written ones, and overwhelmingly prefer them over the reference summaries traditionally used to fine-tune language models (Goyal et al., 2022; Pu et al., 2023; Zhang et al., 2024). Recent work focuses on refining the summarization prompts, e.g., through Chain-of-Thought approaches that first prompt the model to list important facts and then integrate these facts into a summary (Wang et al., 2023) or that guide the model in retrieving relevant context (Liu et al., 2024).
@@ 4 Experiments
We describe our experimental procedure, starting with the dataset and models used [^1] Code and results are available at: https://github.com/enricoliscio/moral-summarization .
@@ 4.1 Dataset
We experiment with the EMONA dataset (Lei et al., 2024), a collection of 400 news articles spanning different topics and political biases. In each sentence of the articles, all event instances were annotated by humans—where an event refers to an occurrence or action, considered the basic element in storytelling (Zhang et al., 2021)—yielding a total of over 10k sentences and 45k events. Each event was then labeled as moral-laden or neutral, with 21% (9.6k) of the events labeled as moral-laden. When labeled as moral-laden, the event was then annotated with one of the ten elements of the Moral Foundation Theory (such as loyalty or fairness). Examples and additional details are provided in Appendix D.3.1.
@@ 4.2 Models
As summarizers, we compare three state-of-the-art models that demonstrate strong general knowledge understanding and are instruction-tuned, as recommended by Zhang et al. (2024)—specifically, Llama-3-70B, Command R+, and DeepSeek R1.
For the Class method, we train Llama-3-8B to identify moral-laden words in the article text in a supervised fashion using the EMONA dataset (we employ a smaller model than the summarization models to keep training and validation manageable). We feed the article’s sentences individually to the classifier which predicts moral-ladenness at a token level. We also tested two other approaches, namely (1) using the full article as input, and (2) a sequential approach that first identifies the sentences that contain moral-laden words and then the moral-laden words in them, but found them to perform worse. Hyperparameters and additional details are reported in Appendix B.
@@ 4.3 Summaries generation procedure
To obtain the classifier for the Class method, we (1) perform a 3-fold cross-validation dividing the training set further into training and validation sets and choose the hyperparameters that lead to the best performance on the validation set, (2) train the model on the whole training set, and (3) perform inference on the test set’s articles. We then generate summaries of the test set articles with the five methods and three summarization models. To account for variability, we repeat the generation five times with fixed seeds and report the average results. We randomly choose one of the five seeds and perform human evaluation on the summaries generated with that seed. For the word-preserving approaches, we treat all annotated moral-laden events as human-annotated moral-laden words, and all other article words as neutral. We do not distinguish among the ten MFT elements in the summary generation, but we do in the evaluation (by measuring moral divergence, see Section 4.4.1).
@@ 4.4 Evaluation
We perform automated, crowd, and expert evaluations of the summaries. Our evaluations are not intended to judge the overall quality of LLM-generated summaries, which has already been extensively studied (Goyal et al., 2022; Pu et al., 2023; Zhang et al., 2024). Instead, we employ these evaluations to compare moral framing preservation across methods and models, and assess whether instructing the summarizer to preserve moral-laden words compromises the overall quality of the summary. First, we employ five automated evaluation metrics to compare the generated summaries across summarization models. We use these results to choose the best-performing summarization model. Then, we perform crowd and expert evaluations of the summaries generated by the chosen model on the test set to investigate the aspects of moral framing preservation that are not captured by the automated metrics. Appendix C provides additional details on the human evaluations.
@@ 4.4.1 Automated evaluation
In line with a recent benchmarking survey (Zhang et al., 2024), we employ the following reference-free automated evaluation metrics to measure the quality of the summaries: (1) SummaC(Laban et al., 2022), which measures how the summary sentences are entailed from the news article; (2) QAFactEval(Fabbri et al., 2022), which measures entailment based on questions generated from the article and answers based on the summary; and (3) BLANC(Vasilyev et al., 2020), which measures the performance boost gained by a language model with access to the summary while performing a language understanding task on the article.
In addition, we introduce the following metrics to measure the preservation of the moral-laden words in the generated summaries: (1) MoralCount, which reports the number of words that were annotated as moral-laden in the article and that are preserved in the summary (to avoid double-counting, we employ the lemmas of the annotated and preserved words). (2) MoralDivergence, which measures the divergence between the distribution of moral annotations in the article and the summary. We normalize the distribution of moral annotations across the ten MFT moral categories and employ the Jensen-Shannon divergence to calculate the divergence between the distribution in the article and the summary (smaller divergence indicates more similarity between the distributions).
@@ 5.1 Automated evaluation
Summary Length Llama-3-70B 153.0 161.4 152.4 174.6 173.9
Command R+ 173.8 191.7 201.3 197.1 194.7
DeepSeek R1 146.8 162.2 159.2 174.3 173.2
Table 2: Results of the automated evaluation of the summaries of the test set articles. As detailed in Section 4.4.1, we report three reference-free automated evaluation metrics, two novel metrics to measure the preservation of moral-laden words, and the average length of the summaries. $\uparrow$ indicates that higher scores are better, $\downarrow$ indicates the opposite.
Second, on MoralCount and MoralDivergence, the methods rank consistently across models in the order: Plain, Direct, CoT, Class, Oracle. Thus, instructing the summarizer to preserve moral framing even without providing a list of moral-laden words (Direct) leads to better preservation of moral-laden words than just instructing to generate a summary (Plain).
@@
Third, as expected, the word-preserving methods outperform the basic methods on MoralCount and MoralDivergence. To compare the three methods, we first report the word-level moral-ladenness prediction performance. Recall (Table 1) that, in CoT, the summarizer was prompted to identify the moral-laden words to be preserved in the summary; in Class, a classifier was trained to detect moral-laden words in the article. The efficacy of these components is instrumental in the subsequent summary generation with the two methods. Table 3 reports their performance, measured by comparing the lists of predicted words and the list of words annotated as moral-laden in the EMONA dataset (thus, random prediction would lead to a near-zero $F_{1}$-score). We observe that Class leads to better moral-ladenness prediction results than CoT, which is reflected in the MoralCount and MoralDivergence results in Table 2. Unsurprisingly, Oracle—where the summarizer is provided with the list of words annotated as moral-laden—outperforms the other two word-preserving methods.
Finally, we observe that the requested maximum summary length (200 words) is generally respected but with variations across methods and models (with the average article length being 738.0 words). Command R+ generates consistently longer summaries than Llama-3-70B and DeepSeek R1, in several cases exceeding the word limit. We observe similar trends for Llama-3-70B and DeepSeek R1, with Oracle and Class leading to the longest summaries, likely due to the efforts in preserving the identified moral-laden words. However, despite the instructions to preserve moral-laden words, CoT results in shorter summaries than the other two word-preserving approaches. This is likely due to the number of moral-laden words that the summarizer is tasked to preserve in the summary—the median number of such words is 11 for CoT, 16.5 for Class, and 21.5 for Oracle. Not being tasked with these instructions, Plain and Direct lead to shorter summaries.
@@ 5.2 Crowd evaluation
Table 4 reports, for each method, the average Likert scores the workers gave to the morally framed spans they highlighted, and the differences between the scores assigned to the summaries generated by different methods.
@@
First, we observe that Class and Oracle emerge as the best-performing methods with significant differences from the other three methods (as shown by the Wilcoxon pairwise comparison matrix in Table 4). However, we notice no significant difference between the two, although Oracle is more effective in preserving the annotated moral-laden words than Class, which in turn is more effective than CoT (as shown in Section 5.1). These results suggest that, when a sufficient moral-ladenness prediction performance is met, preserving moral-laden words in the summary improves the preservation of moral framing, but a better moral-ladenness prediction performance does not necessarily correlate with better moral framing preservation.
Second, we observe that all average scores range between 3 and 3.5 (rightmost column of Table 4). This is not surprising since these scores evaluate the extent to which morally framed article spans are preserved in the summary, and some are inevitably lost due to the summary’s shorter length.
Third, we find no significant correlations between automated and crowd evaluation results (as reported in Table 5.2 together with the correlation to the summaries length). Thus, judging moral framing preservation requires human evaluation, as it (1) is not correlated to the summary quality, and (2) cannot be assessed by measuring the preservation of moral words.
Finally, we note that traditional inter-annotator agreement metrics are not applicable in our setting. For each article, annotators highlighted morally framed spans and rated each one on a Likert scale; these spans often differed significantly between annotators in both number and content. In addition, crucially, aligning on span selection was not the goal; highlighting served as a mnemonic tool to recall morally framed content. Our focus is on how well annotators believe these moral frames are preserved in the summaries, as reflected in their Likert ratings. We explore a possible way to assess rating consistency in Appendix D.2.
@@
To evaluate the effect of preserving moral-laden words in the summary, we examine the scores given to the highlighted spans containing at least one annotated moral-laden word that is preserved in the summary. The spans containing moral-laden words are the same for all summaries—however, some summaries preserve these words while others do not. Thus, the number of these spans varies across methods. Table 6 reports the average scores assigned to such spans and their number. We observe that the scores are similar across methods and consistently higher than the average scores presented in Table 4, suggesting that maintaining moral-laden words promotes the preservation of moral framing. The difference across methods lies in the number of such highlighted spans, with Class and Oracle counting more of such spans (respectively, 24% and 39% more when compared to Plain), leading to the difference in the average scores observed in Table 4.
@@
The workers were allowed to highlight as many morally framed spans as necessary, resulting in a number of highlights ranging from 0 (in a single case) to 31, with an average of $6.6\pm 4.7$. We correlate the number of spans a worker highlighted to the Likert scores they assigned to the summaries. We group the number of highlights per article in blocks of three and report in Figure 2 the average scores assigned by the workers to each method’s summaries. The scores exhibit a downward trend—the more spans a worker highlighted, the lower the average scores they assigned to them. This is expected since the more text spans are highlighted, the more challenging it is for the summarizer to preserve them all in the summary. Next, we observe that Class generally outperforms the other methods across the numbers of highlights. In contrast, Oracle leads to the worst results in the 1-2-3 range but the best in the 4-5-6 range. We conjecture that this is due to the number of moral-laden words to be preserved in the summary (which is the highest for Oracle, as reported in Section 5.1)—attempting to preserve a large number of words may dilute the reporting of key aspects of the article, which is especially noticeable when only a few key morally framed text spans are highlighted and evaluated.
[…]
@@ 6 Conclusions
Explaining how the summarization models identify moral-laden words and decide which ones to preserve is instrumental to improving transparency to users, especially considering the inherent subjectivity of morality (van der Meer et al., 2024). Reducing reliance on human annotations can foster the approach’s practicality, e.g., by enhancing the Chain-of-Thought method with the support of few-shot examples or moral lexicons, or by evaluating how LLM judges can approximate human judgements on this task. Finally, the EMONA dataset offers further avenues of exploration, e.g., our results could be correlated with the political bias of the article’s source, and the prompting methods could be extended by integrating the Moral Foundation Theory annotations associated with moral-laden words.
Acknowledgments
This material is produced as part of AlgoSoc, a collaborative 10-year research program on public values in the algorithmic society, and the Hybrid Intelligence Centre, both funded by the Dutch Ministry of Education, Culture and Science (OCW) under the Gravitation programme (project numbers 024.005.017 and 024.004.022). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of OCW or those of the AlgoSoc consortium as a whole. Research reported in this work was partially or completely facilitated by computational resources and support of the Delft AI Cluster (DAIC) at TU Delft (RRID: SCR_025091), but remains the sole responsibility of the authors, not the DAIC team. Michela Lorandi’s work was conducted with the financial support of the Science Foundation Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real) under Grant No 18/CRT/6224 and the ADAPT SFI Research Centre at Dublin City University, funded by the Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2. A special thanks to Mohammed Al Owayyed for his support with setting up the crowd task.
@@ Acknowledgments
We provide all the necessary details for reproducing our experiments throughout the paper and the Appendix. Methods, prompting, and training procedures are described in Sections 3 and 4. Appendix A reports the complete prompts used to generate the summaries. Appendix B provides additional details on our experiments, including the tested and used hyperparameters and computational infrastructure. Appendix C presents additional information on the human evaluations, including the aggregated demographics and further details on the crowd task. Code, generated summaries, and (anonymized) automated, crowd, and expert evaluation results are available on GitHub1. The model weights trained with the Class method are available in a public repository (Liscio et al., 2025a).